OpenStudy (lovelyanna):

Evaluate the expression when a = –1.4 and b = –2.7. |a| – b A. 4.1 B. 1.3 C. –1.3 D. –4.1

OpenStudy (the_fizicx99):

4.1*

OpenStudy (lovelyanna):

Thanks, can you please explain?

OpenStudy (the_fizicx99):

Well because its |a| – b I would add -1.4 and -2.7 which is -4.1 and a negative + negative = positive

OpenStudy (lovelyanna):

Oh, but I think it's -1/3, or 1/3?

OpenStudy (the_fizicx99):

-1/3 ?

OpenStudy (lovelyanna):

This question is confusing.. :/

OpenStudy (the_fizicx99):

If you're referring to doing 41/10 = 4 + 1/10 soo?

OpenStudy (lovelyanna):

.....??

OpenStudy (the_fizicx99):

41/10 = 4.1 1/10 = .1 + 4 = 4.1